10.0 mg/dl calibration reference had an absorbance of 628 bance of .270. The 24 example: A urinary calcium level (mg Cat+ /24 hours) for the following while a specimen urine diluted 1:2 gave an absor hour uri ne volume was 1320 ml. (Show all work) Is this value normal or abnormal? Explain. (6 points)
Answer #1

Calculation of 24 Hour urine calcium by photometry:

Ca2+ (mg/dl) = (Aunknown/Astandard) x 10

Aunknown = Absorbance of given urine sample after dilution with the reagent

Astandard = Absorbance of urine sample with 10 mg/dl calibration reference

Aunknown = 0.270

Astandard = 0.628

Ca2+ (mg/dl) = (0.270/0.628) x 10

                      = 4.299 mg/dl

Volume of 24 hour urine sample is 1320 ml or 13.2 dl

Ca2+ in 24 hour Urine sample 4.299 x 13.2 = 56.74 mg/day

Now, the calcium level in urine of a person with normal diet ranges from 100 – 300 mg/day. A value of 56.74 mg/day indicates low calcium diet. Other causes of low urine calcium can be malabsorption syndrome, hypoparathyroidism, Vitamin D deficiency and use of Thiazide diuretics.